OpenStudy (nsh4267):

Solve the equation (f o g)(x) = 0. f(x)=x^2-3 g(x)= x+5 x=?

OpenStudy (anonymous):

x=-5+square root(3) and -5-square root(3)

OpenStudy (anonymous):

(fog)(x)=f(g(x))=f(x+5)=(x+5)^2-3 =0 solving this we get the answer

OpenStudy (anonymous):

@nsh4267

OpenStudy (anonymous):

f(g(x))=(x+5)^2-3 0=(x+5)^2-3 3 = (x+5)^2 +-sqrt 3 = x+5 x=-5+- sqrt3